# Cost-Sensitive Learning to Defer to Multiple Experts with Workload Constraints

### v2

@@ 3.1 Deferral Formulation
To obtain $h^{*}$, we can train a binary classifier using any proper binary composite loss with a well defined inverse link function $\psi^{-1}$, such as the logistic loss [Reid and Williamson, 2010], for which $\psi^{-1}(g)=1/(1+e^{-g})$. The empirical estimator of the expected value of the logistic loss is given by
$\begin{split}\mathcal{L}_{\log}(g)&=\frac{1}{N}\sum_{i=1}^{N}\bigg{[}-\mbox{y}_{i}\log\bigg{(}\psi^{-1}\big{(}g(\bm{x}_{i})\big{)}\bigg{)}-(1-\mbox{y}_{i})\log\bigg{(}1-\psi^{-1}\big{(}g(\bm{x}_{i})\big{)}\bigg{)}\bigg{]},\end{split}$ (3)
where $g(\bm{x}_{i})$ is the score of sample $\bm{x}_{i}$. Proper binary composite losses ensure that $\mathbb{E}[\mathcal{L}]$ is minimized when $\psi^{-1}(gg(\bm{x}))=\mathbb{P}(\mbox{y}=1\lvert\bm{x})$ [Buja et al., 2005, Reid and Williamson, 2010], meaning that minimization of this empirical loss will converge to $g^{*}$ as $N\xrightarrow{}\infty$, a scoring function producing calibrated estimates of $P(\mbox{y}=1\lvert\bm{x})$. As such, defining
$h^{*}(\bm{x})=\begin{cases}1\;\mbox{if}\;\psi^{-1}\big{(}g^{*}(\bm{x})\big{)}>0.5\\ 0\quad\mbox{otherwise},\end{cases}$ (4)
ensures that $h^{*}$, in the large sample limit, will agree with the Bayes-optimal classifier $h^{B}$.
@@
In order to obtain $r^{*}$, we resort to a human expertise model (HEM) with scoring function $g_{\perp}:\mathcal{X}\cup\{1,...,J\}\rightarrow\mathbb{R}$ , predicting whether an expert is correct (1) or incorrect (0). By conditioning the model’s output on the expert index, we can train this model using the training set $S$, which only contains one expert prediction per instance. Although it would be possible to train one scoring function per expert on the subset of the data with said expert’s predictions (which would be similar to the approach of Verma et al. [2023]), we choose to use a single model for the entire team. In scenarios with limited data pertaining to each expert, it can be beneficial to model the team’s behavior jointly, as there may not be enough data to train reliable models for every single expert. By conditioning the model’s output on the expert’s index, we still allow the model to adapt to each individual expert’s decision-making processes, should that be beneficial. Note that, should additional information $\bm{e}\in\mathcal{E}$ be available to the experts (e.g. ML model score), the HEM scoring function can take the form $g_{\perp}:\mathcal{X}\cup\{1,...,J\}\cup\mathcal{E}\rightarrow\mathbb{R}$, thus conditioning the expert correctness probability estimate on all the information available to the expert.
We obtain the optimal scoring function $g_{\perp}^{*}$ by minimizing the logistic loss over the training set $S$, which allows obtaining the estimates of $\mathbb{P}(\mbox{y}=m_{j}\lvert\bm{x},j)$, given by $\psi^{-1}\big{(}g^{*}_{\perp}(\bm{x},j)\big{)}$. As such, by defining the rejector $r^{*}$ as
$r^{*}(\bm{x})=\begin{cases}0\;\mbox{if}\;\max\{\psi^{-1}\big{(}g^{*}(\bm{x})\big{)},1-\psi^{-1}\big{(}g^{*}(\bm{x})\big{)}\}>\psi^{-1}\big{(}g^{*}_{\perp}(\bm{x},j)\big{)}\quad\forall j\in\{1,...,J\}\\ \operatorname*{arg\,max}_{j\in\{1,...,J\}}\psi^{-1}\big{(}g^{*}_{\perp}(\bm{x},j)\big{)}\quad\mbox{otherwise,}\end{cases}$ (5)
then $r^{*}$ will agree with the Bayes-optimal rejector $r^{B}$, thus proving that our approach in modeling and training $r^{*}$ and $h^{*}$ under limited data availability will yield a classifier-rejector pair that converges to the minimizers of $L_{0-1}$, $h^{B}$, and $r^{B}$.
@@ 3.2 Definition of Capacity Constraints
To shift focus to deferral under capacity constraints, we start by formalizing the work capacity limitations of the expert team. Humans are limited in the number of instances they may process in any given time period (e.g., work day). In real-world systems, human capacity must be applied over batches of instances, not over the whole dataset at once (e.g., balancing the human workload over an entire month is not the same as balancing it daily). A real-world assignment system must then process instances taking into account the human limitations over a given “batch” of cases, corresponding to a pre-defined time period. We divide our dataset into several batches and, for each batch, define the maximum number of instances that can be processed by each expert. In any given dataset comprised of $N$ instances, capacity constraints can be represented by a vector $\bm{b}$, where component $b_{i}$ denotes which batch the instance $i\in\{1,...,N\}$ belongs to, as well as a human capacity matrix $\bm{H}$, where element $H_{b,j}$ is a non-negative integer denoting the number of instances in batch $b$ the human expert $j$ can process.
@@ 3.3 Global Loss Minimization under Capacity Constraints
In the previous sections, we detailed our approach to train the classifier-rejector pair, showing that $h^{*}$ and $r^{*}$ converge to the Bayes-optimal $h^{B}$ and $r^{B}$. Note that this classifier-rejector formulation produces optimal point-wise solutions, disregarding any work capacity limitations. Should an expert be more likely to be correct than all their colleagues and the ML classifier over the entire feature space, the set of optimal point-wise assignments would be to always defer instances to said expert, a clearly unfeasible solution in any real-world system.
@@
A point-wise formulation of $h^{*}$ and $r^{*}$ is clearly inadequate to optimize the assignments over a set of instances while respecting expert capacity constraints. However, their scoring functions $g^{*}$ and $g_{\perp}^{*}$ will still be of use in providing calibrated estimates of the probability of correctness. To formulate the assignment problem under capacity constraints, we consider that our objective is to maximize the expected probability of correctness over all instances to be deferred. Recall that $\mathcal{Y}=\{0,1\}$ and consider the assigner $a:\mathcal{X}\rightarrow\mathcal{Y}\cup\{-1,...,-J\}$, where $a(\bm{x})=-j$ denotes the decision to defer instance $\bm{x}$ to the $j$th expert, whereas $a(x)=y$, with $y\in\mathcal{Y}$, is an automatic prediction of class $y$. The estimate of the probability of correctness for all possible assignments on a given instance is given by
@@
^$\begin{split}\hat{\mathbb{P}}(\mbox{correct}\lvert\bm{x})=\begin{cases}1-\psi^{-1}\big{(}g^{*}(\bm{x})\big{)}\;\mbox{if}\;a(\bm{x})=0\\ \psi^{-1}\big{(}g^{*}(\bm{x})\big{)}\;\mbox{if}\;a(\bm{x})=1\\ \psi^{-1}\big{(}g^{*}_{\perp}(\bm{x},j)\big{)}\;\mbox{if}\;a(\bm{x})=-j.\end{cases}\end{split}$ (6)
@@
To represent the assignment decisions over a batch $b$ comprised of $n_{B}$ instances, without loss of generality indexed in $\{1,...,n_{B}\}$, consider the $n_{B}\times J$ matrix of assignments $\bm{A}$, where each element $A_{i,j}$ is a binary variable that denotes if instance $i$ is deferred to expert $j\in\{1,...,J\}$. The optimal set of assignments is given by
@@
$\begin{split}\bm{A}^{*}&=\operatorname*{arg\,min}_{A\in\{0,1\}^{n_{B}\times J}}\sum_{i=1}^{n_{B}}\sum_{j=1}^{J}\hat{\mathbb{P}}(\mbox{correct}\lvert\bm{x}_{i})A_{i,j},\\ \text{s.t.}\quad\sum_{i=1}^{n_{B}}A_{i,j}&=H_{b,j},\;\mbox{for}\;j\in\{1,2,...,J\}\;\mbox{and}\;\sum_{j=1}^{J}A_{i,j}=1,\;\mbox{for}\;i\in\{1,2,...,n_{B}\}.\end{split}$ (7)
The first constraint refers to human decision capacity: with an equality constraint, the number of instances assigned to each decision-maker is predefined in the problem statement; this constraint may be changed to an inequality expressing the maximum number of assignments per decision-maker. The second constraint states that each instance must be assigned to one and only one decision-maker. We solve the assignment problem 7 using the constraint programming solver CP-SAT from Google Research’s OR-Tools [Perron and Didier, 2023].
@@ 3.4 Cost-Sensitive Learning
Finally, having discussed how to optimize for the $0-1$ loss under capacity constraints, we now focus on adapting our method to work under an arbitrary cost structure. Training a classifier $k$ with score function $g_{k}$ involves approximating the expected value of some surrogate loss function $\mathcal{L}$ by the empirical average of the point-wise losses over the training set:
$\mathbb{E}_{(\bm{x},y)\sim D}[\mathcal{L}(\bm{x},g_{k}(\bm{x}))]\approx\frac{1}{N}\sum_{i}^{N}\mathcal{L}(\bm{x}_{i},g_{k}(\bm{x}_{i})),,$ (8)
@@
where $D$ denotes the data distriution. However, instances may have different misclassification costs, in which case a surrogate for the 0-1 loss is not adequate. Assuming that each instance $\bm{x}_{i}$ has an associated misclassification cost $c_{i}$, the goal is then to learn a classifier that minimizes the expected cost, $\mathbb{E}_{\bm{x},y,c\sim D}[c\;\mathbb{I}(k(\bm{x})\neq y)]$. Minimizing surrogates to the 0-1 loss ensures that we are minimizing $\mathbb{E}_{\bm{x},y,c\sim D}[\mathbb{I}(k(\bm{x})\neq y)]$, which is misaligned with our objective. Zadrozny et al. [2003] show that if we have examples drawn from a different distribution
$\begin{split}&\tilde{D}(\bm{x},y,c)=\frac{c}{\mathbb{E}_{c\sim D}[c]}D(\bm{x},y,c),\;\mbox{then}\;\mathbb{E}_{\bm{x},y,c\sim\tilde{D}}[\mathbb{I}(k(\bm{x})\neq y)]=\frac{1}{\mathbb{E}_{c\sim D}[c]}\mathbb{E}_{\bm{x},y,c\sim D}[c\,\mathbb{I}(k(\bm{x})\neq y)].\end{split}$ (9)
@@
Equation 9 shows that selecting the decision rule $k$ to minimize the error rate under $\tilde{D}$ is equivalent to selecting $k$ to minimize the expected misclassification cost under $D$. This means we can obtain our desired classifier by training it under the distribution $\tilde{D}$, using the softmax log-loss. To train a classifier under $\tilde{D}$, a common approach [Zadrozny et al., 2003, Elkan, 2001] is to re-weight the instances according to their costs. In this way, we obtain a classifier $k$ which prioritizes correctness according to the instances’ weights, by minimizing the empirical estimation of the misclassification cost:
$\frac{1}{N\,\mathbb{E}_{c\sim D}[c]}\sum_{i=1}^{N}c_{i}\mathcal{L}(\bm{x}_{i},g_{k}(\bm{x}_{i})).$ (10)
As such, to adapt our method to cost sensitive scenarios, we simply reweight the logistic loss when training the scoring functions $g$ and $g_{\perp}$.

### v3

@@ 3.1 Deferral Formulation
To obtain $\hat{h}$, we can train a binary classifier using any proper binary composite loss with a well defined inverse link function $\psi^{-1}$, such as the logistic loss (Reid & Williamson, 2010), for which $\psi^{-1}(g)=1/(1+e^{-g})$. The empirical estimator of the expected value of the logistic loss is given by
$\begin{split}\mathcal{L}_{\text{classifier}}\big{(}\{\bm{x}_{i},y_{i},c_{i}\}_{i=1}^{N},g\big{)}&=\frac{1}{N}\sum_{i=1}^{N}\bigg{[}-y_{i}\log\bigg{(}\psi^{-1}\big{(}g(\bm{x}_{i})\big{)}\bigg{)}-(1-y_{i})\log\bigg{(}1-\psi^{-1}\big{(}g(\bm{x}_{i})\big{)}\bigg{)}\bigg{]},\end{split}$ (4)
where $g(\bm{x}_{i})$ is the score of sample $\bm{x}_{i}$. Proper binary composite losses ensure that $\mathbb{E}[\mathcal{L}]$ is minimized when $\psi^{-1}(g(\bm{x}))=\mathbb{P}(y=1\lvert\bm{x})$ (Buja et al., 2005; Reid & Williamson, 2010), meaning that minimization of this empirical loss will converge to $\hat{g}$ as $N\xrightarrow{}\infty$, a scoring function producing calibrated estimates of $P(y=1\lvert\bm{x})$. As such, defining
$\hat{h}(\bm{x})=\begin{cases}1\;&\mbox{if}\;\psi^{-1}\big{(}\hat{g}(\bm{x})\big{)}>0.5\\ 0\;&\mbox{otherwise},\end{cases}$ (5)
ensures that $\hat{h}$, in the large sample limit, will agree with the Bayes-optimal classifier $h^{*}$.
@@
In order to obtain $\hat{r}$, we resort to a human expertise model (HEM) with scoring function $g_{\perp}:\mathcal{X}\times\{1,...,J\}\rightarrow\mathbb{R}$ , predicting whether an expert is correct (1) or incorrect (0). By conditioning the model’s output on the expert index, we can train this model using the training set $S$, which only contains one expert prediction per instance. Although it would be possible to train one scoring function per expert on the subset of the data with said expert’s predictions (which would be similar to the approach of Verma et al. (2023)), we choose to use a single model for the entire team. In scenarios with limited data pertaining to each expert, it can be beneficial to model the team’s behavior jointly, as there may not be enough data to train reliable models for every single expert. By conditioning the model’s output on the expert’s index, we still allow the model to adapt to each individual expert’s decision-making processes, should that be beneficial. Note that, should additional information $\bm{e}\in\mathcal{E}$ be available to the experts (e.g., ML model score), the HEM scoring function can take the form $g_{\perp}:\mathcal{X}\cup\{1,...,J\}\cup\mathcal{E}\rightarrow\mathbb{R}$, thus conditioning the expert correctness probability estimate on all the information available to the expert. We obtain the optimal scoring function $\hat{g}_{\perp}$ by minimizing
$\begin{split}\mathcal{L}_{\text{HEM}}\big{(}\{\bm{x}_{i},y_{i},m_{j,i},j_{i}\}_{i=1}^{N},g_{\perp}\big{)}=\frac{1}{N}\sum_{i=1}^{N}\bigg{[}&-\mathbb{I}[m_{j,i}=y]\log\Big{(}\psi^{-1}\big{(}g_{\perp}(\bm{x}_{i},j)\big{)}\Big{)}\\ &-\mathbb{I}[m_{j,i}\neq y]\log\Big{(}1-\psi^{-1}\big{(}g_{\perp}(\bm{x}_{i},j)\big{)}\Big{)}\bigg{]},\end{split}$ (6)
where $g_{\perp}(\bm{x}_{i},j)$ is the score associated with deferring instance $\bm{x}_{i}$ to expert $j$. This allows us to obtain the estimates of $\mathbb{P}(y=m_{j}\lvert\bm{x},j)$, given by $\psi^{-1}\big{(}\hat{g}_{\perp}(\bm{x},j)\big{)}$. As such, defining the rejector $\hat{r}$ as
$\hat{r}(\bm{x})=\begin{cases}0\;&\mbox{if}\;\max\{\psi^{-1}\big{(}\hat{g}(\bm{x})\big{)},1-\psi^{-1}\big{(}\hat{g}(\bm{x})\big{)}\}>\psi^{-1}\big{(}\hat{g}_{\perp}(\bm{x},j)\big{)},\;\forall j\in\{1,...,J\}\\ \displaystyle\operatorname*{arg\,max}_{j\in\{1,...,J\}}\psi^{-1}\big{(}\hat{g}_{\perp}(\bm{x},j)\big{)}\;&\mbox{otherwise,}\end{cases}$ (7)
implies that $\hat{r}$ will agree with the Bayes-optimal rejector $r^{*}$, thus proving that our approach in modeling and training $\hat{r}$ and $\hat{h}$ under limited data availability yields a classifier-rejector pair that converges to $h^{*}$ and $r^{*}$, the minimizers of $L_{0-1}$,
@@ 3.2 Definition of Capacity Constraints
To shift focus to deferral under capacity constraints, we start by formalizing the work-capacity limitations of the expert team. Humans are limited in the number of instances they may process in any given time period (e.g., work day). In real-world systems, human capacity must be applied over batches of instances, not over the whole dataset at once (e.g., balancing the human workload over an entire month is not the same as balancing it daily). A real-world assignment system must then process instances taking into account the human limitations over a given “batch” of cases, corresponding to a pre-defined time period. We divide our dataset into several batches and, for each batch, define the maximum number of instances that can be processed by each expert. In any given dataset comprised of $N$ instances, divided into $n_{\text{batches}}$, capacity constraints can be represented by a vector $\bm{b}$, where component $b_{i}$ denotes which batch the instance $i\in\{1,...,N\}$ belongs to, as well as a human capacity matrix $\bm{H}\in\mathbb{N}_{0}^{n_{\text{batches}}\times J}$, where element $H_{b,j}$ is a non-negative integer denoting the number of instances in batch $b$ that human expert $j$ can process.
@@ 3.3 Global Loss Minimization under Capacity Constraints
In the previous sections, we detailed our approach to train the classifier-rejector pair, showing that $\hat{h}$ and $\hat{r}$ converge to the Bayes-optimal $h^{*}$ and $r^{*}$. Note that this classifier-rejector formulation produces optimal point-wise solutions, disregarding any work-capacity limitations. Should an expert be more likely to be correct than all their colleagues and the ML classifier over the entire feature space, the set of optimal point-wise assignments would be to always defer instances to said expert, a clearly unfeasible solution in any real-world system.
@@
A point-wise formulation of $\hat{h}$ and $\hat{r}$ is clearly inadequate to optimize the assignments over a set of instances while respecting expert capacity constraints. However, the scoring functions $\hat{g}$ and $\hat{g}_{\perp}$, obtained in the training steps detailed above, will still be of use in providing calibrated estimates of the probability of correctness. To formulate the assignment problem under capacity constraints, we consider that our objective is to maximize the expected probability of correctness over all instances to be deferred. Recall that $\mathcal{Y}=\{0,1\}$ and consider the assignment decision for each instance $a_{i}\in\{1,...,J+2\}$, where $a_{i}=y+1$, with $y\in\mathcal{Y}$, is an automatic prediction of class $y$ for instance $i$, whereas $a_{i}=j+2$ denotes the decision to defer instance $i$ to the $j$th expert. The estimate of the probability of correctness for all possible assignments on a given instance is then given by
@@
^$\begin{split}\hat{\mathbb{P}}(\mbox{correct}\lvert\bm{x}_{i},a_{i})=\begin{cases}1-\psi^{-1}\big{(}\hat{g}(\bm{x}_{i})\big{)}\;&\mbox{if}\;a_{i}=1\\ \psi^{-1}\big{(}\hat{g}(\bm{x}_{i})\big{)}\;&\mbox{if}\;a_{i}=2\\ \psi^{-1}\big{(}\hat{g}_{\perp}(\bm{x}_{i},j)\big{)}\;&\mbox{if}\;a_{i}=j+2.\end{cases}\end{split}$ (8)
@@
To represent the assignment decisions over a batch $b$ comprised of $n_{B}$ instances, without loss of generality indexed in $\{1,...,n_{B}\}$, consider the $n_{B}\times(2+J)$ matrix of assignments $\bm{A}$, where each element $A_{i,a_{i}}$ is a binary variable that denotes if the assignment decision $a_{i}$ is taken for instance $i$. The optimal set of assignments is given by
@@
$\begin{split}\bm{A}^{*}=\operatorname*{arg\,min}_{A\in\{0,1\}^{n_{B}\times(2+J)}}&\sum_{i=1}^{n_{B}}\sum_{a_{i}=1}^{J+2}\hat{\mathbb{P}}(\mbox{correct}\lvert\bm{x}_{i},a_{i})A_{i,a_{i}},\\ \mbox{s.t.}&\;\,\sum_{i=1}^{n_{B}}A_{i,a_{i}}=H_{b,a_{i}},\;\mbox{for}\;a_{i}\in\{3,...,J+2\},\;\\ &\;\sum_{a_{i}=1}^{J+2}A_{i,a_{i}}=1,\;\mbox{for}\;i\in\{1,2,...,n_{B}\}.\end{split}$ (9)
The first constraint refers to human capacity: with an equality constraint, the number of instances assigned to each decision-maker is predefined in the problem statement; this constraint may be changed to an inequality expressing the maximum number of assignments per decision-maker. The second constraint states that each instance must be assigned to one and only one decision-maker. We solve the assignment problem 9 using the constraint programming solver CP-SAT from Google Research’s OR-Tools (Perron & Didier, 2023). Finding solutions to these combinatorial problems is not trivial, as solution times grow exponentially with the problem’s size (Huberman et al., 1997). CP-SAT is a parallel portfolio solver, leveraging the available processor cores to run different solution search strategies on each. By doing so, portfolio solvers aim to more efficiently cover the solution-space, under the assumption that different algorithms, using different heuristics, may compliment each other’s weaknesses (Huberman et al., 1997; Gomes & Selman, 2001). This approach proves effective as selecting the optimal search algorithm for a given constraint problem is considered a non-trivial task. We selected CP-SAT in our implementation as it has been repeatedly shown to be the best-performing publicly available solver in a wide array of constraint programming problems in the MiniZinc challenge (Stuckey et al., 2014), where solvers are tasked to finding the optimal or a near-optimal solution within a time-limit. While this solver can find feasible solutions relatively quickly, proving a solution’s optimality can take an extremely long time, meaning that in practice it is often better to set a time limit, having the system return the best solution it found so far. On our machine (Intel(R) Xeon(R) Gold 5120 CPU @ 2.20GHz), no significant improvements were obtained for timeout limits above 60 seconds.
@@ 3.4 Cost-Sensitive Learning
Finally, having discussed how to optimize for the 0-1 loss under capacity constraints, we now focus on adapting our method to work under an arbitrary cost structure. To do so, we follow an instance re-weighting approach (Zadrozny et al., 2003; Elkan, 2001), where each point-wise loss over the training set is multiplied by the cost $c_{i}$ associated with said instance. This guarantees that the minimization of the surrogate losses used to train $g$ and $g_{\perp}$ will result in scoring functions that minimize the misclassification cost instead of the error-rate. A more detailed description of the re-weighting approach follows.
@@
Training a classifier $h$ with score function $g$ involves approximating the expected value of its surrogate loss function $\ell$ by the empirical average of the point-wise losses over the training set,
$\mathbb{E}_{(\bm{x},y)\sim D}[\ell(\bm{x},g(\bm{x}),y)]\approx\frac{1}{N}\sum_{i}^{N}\ell(\bm{x}_{i},g(\bm{x}_{i}),y_{i}),$ (10)
@@
where $D$ denotes the data distribution. Note, however, that instances may have different misclassification costs, in which case a surrogate for the 0-1 loss is not adequate. Assuming that each instance $\bm{x}_{i}$ has an associated misclassification cost $c_{i}$, the goal is then to learn a classifier that minimizes the expected cost, $\mathbb{E}_{\bm{x},y,c\sim D}[c\;\mathbb{I}(h(\bm{x})\neq y)]$. Minimizing surrogates to the 0-1 loss ensures that we minimize $\mathbb{E}_{\bm{x},y,c\sim D}[\mathbb{I}(h(\bm{x})\neq y)]$, which is misaligned with our objective. Zadrozny et al. (2003) show that if we have examples drawn from a different distribution,
$\begin{split}&\tilde{D}(\bm{x},y,c)=\frac{c}{\mathbb{E}_{c\sim D}[c]}D(\bm{x},y,c),\end{split}$ (11)
then
$\begin{split}\mathbb{E}_{\bm{x},y,c\sim\tilde{D}}[\mathbb{I}(h(\bm{x})\neq y)]=\frac{1}{\mathbb{E}_{c\sim D}[c]}\mathbb{E}_{\bm{x},y,c\sim D}[c\,\mathbb{I}(h(\bm{x})\neq y)].\end{split}$ (12)
@@
Equation 12 shows that selecting the decision rule $h$ to minimize the error rate under $\tilde{D}$ is equivalent to selecting $h$ to minimize the expected misclassification cost under $D$. This means we can obtain our desired classifier by training it under the distribution $\tilde{D}$, using the log-loss. To train a classifier under $\tilde{D}$, a common approach (Zadrozny et al., 2003; Elkan, 2001) is to re-weight the instances according to their costs. In this way, we obtain a classifier $\hat{h}$ that prioritizes correctness according to the instances’ weights, by minimizing the empirical estimate of the misclassification cost:
$\frac{1}{N\,\mathbb{E}_{c\sim D}[c]}\sum_{i=1}^{N}c_{i}\ell(\bm{x}_{i},g(\bm{x}_{i}),y_{i}).$ (13)
As such, we just have to reweight the instances in $\mathcal{L}_{\text{classifier}}$, where $\ell$ is the point-wise log-loss (see Eq. 4), when training the scoring function $g$. We follow the same approach with respect to the rejector, reweighting the instances in $\mathcal{L}_{\text{HEM}}$ (see Eq. 6) when training $g_{\perp}$. The re-weighted empirical losses are given by
$\begin{split}\mathcal{L}^{{}^{\prime}}_{\text{classifier}}\big{(}\{\bm{x}_{i},y_{i},c_{i}\}_{i=1}^{N},g\big{)}&=\frac{1}{N}\sum_{i=1}^{N}c_{i}\bigg{[}-y_{i}\log\Big{(}\psi^{-1}\big{(}g(\bm{x}_{i})\big{)}\Big{)}-(1-y_{i})\log\Big{(}1-\psi^{-1}\big{(}g(\bm{x}_{i})\big{)}\Big{)}\bigg{]},\end{split}$ (14)
$\begin{split}\mathcal{L}^{{}^{\prime}}_{\text{HEM}}\big{(}\{\bm{x}_{i},y_{i},c_{i},m_{j,i},j_{i}\}_{i=1}^{N},g_{\perp}\big{)}=\frac{1}{N}\sum_{i=1}^{N}c_{i}\bigg{[}&-\mathbb{I}[m_{j,i}=y]\log\Big{(}\psi^{-1}\big{(}g_{\perp}(\bm{x}_{i},j)\big{)}\Big{)}\\ &-\mathbb{I}[m_{j,i}\neq y]\log\Big{(}1-\psi^{-1}\big{(}g_{\perp}(\bm{x}_{i},j)\big{)}\Big{)}\bigg{]}.\end{split}$ (15)
The full algorithm is described in the following pseudo-code blocks. Note that, for the training process, the function “Minimize-Loss” represents any optimization algorithm that minimizes an empirical loss (e.g., gradient descent, gradient boosting).
Algorithm 1 Pseudo-Code for DeCCaF - Training
1:Input: Training data $S=\{\bm{x}_{i},y_{i},m_{j,i},c_{i}\}_{i=1}^{N}$
2:Initialize $g$, $g_{\perp}$triangleright Initial scoring functions (e.g., NN with random weights)
3:$\hat{g}\leftarrow\mbox{Minimize-Loss}\Big{(}\mathcal{L}^{{}^{\prime}}_{\text{classifier}}\big{(}\{\bm{x}_{i},y_{i},c_{i}\}_{i=1}^{N},g\big{)}\Big{)}$triangleright see Eq. 14
4:$\hat{g}_{\perp}\leftarrow\mbox{Minimize-Loss}\Big{(}\mathcal{L}^{{}^{\prime}}_{\text{HEM}}\big{(}\{\bm{x}_{i},y_{i},c_{i},m_{j,i}\}_{i=1}^{N},g_{\perp}\big{)}\Big{)}$triangleright see Eq. 15
Algorithm 2 Pseudo-Code for DeCCaF - Inference for Batch $b$
1:Input: Instances $\{\bm{x}_{i}\}_{i=1}^{n_{B}}$, Batch Capacity Constraints $\{\bm{H}_{b,j}\}_{j=i}^{J}$triangleright $\bm{H}\in\mathbb{N}_{0}^{n_{\text{batches}}\times J}$, see Section 3.2
2:Initialize $\bm{P}\in\mathbb{R}^{n_{B}\times(J+2)}$triangleright Stores ^$\hat{\mathbb{P}}(\mbox{correct}\lvert\bm{x}_{i},a_{i})$, see Eq. 8
3:Initialize $\bm{A}\in\{0,1\}^{n_{B}\times(J+2)}$triangleright Stores $A_{i,a_{i}}$ denoting if decision $a_{i}$ is taken for instance i, see Eq. 9
4:for $i\leftarrow\{1,..,n_{B}\}$ do
5:     for $a_{i}\leftarrow\{1,..,J+2\}$ do
6:         if $a_{i}$ = 1 then
7:              $P_{i,a_{i}}\leftarrow 1-\psi^{-1}\big{(}\hat{g}(\bm{x}_{i})\big{)}$
8:         else if $a_{i}$ = 2 then
9:              $P_{i,a_{i}}\leftarrow\psi^{-1}\big{(}\hat{g}(\bm{x}_{i})\big{)}$
10:         else
11:              $P_{i,a_{i}}\leftarrow\psi^{-1}\big{(}\hat{g}_{\perp}(\bm{x}_{i},a_{i}-2)\big{)}$
12:         end if
13:     end for
14:end for
15:$\mathbf{A^{*}}\leftarrow\mbox{CPSolver}\big{(}\bm{P},\{\bm{H}_{b,j}\}_{j=i}^{J}\big{)}$triangleright Optimization Problem in (7)